OpenStudy (anonymous):

Integral of 2sin(2x)cos(2x) dx = Ok... so I used the fact that sin 2x = 2sinxcosx to treat 2x here as x, and so I simplified it to sin 4x If I did everything right. But I am not too sure what to do next... Or what to do with integral of trigs that have functions inside of them. Can anyone point me the way?

OpenStudy (johnweldon1993):

This just looks like a simple u-substitution

OpenStudy (johnweldon1993):

\[\large 2\int^{}_{} sin(2x)cos(2x)dx\]

OpenStudy (johnweldon1993):

Let u = 2x du = 2dx dx = du --- 2

OpenStudy (johnweldon1993):

\[\large \frac{2}{2}\int^{}_{} sin(u)cos(u)du\]

OpenStudy (johnweldon1993):

\[\large \int^{}_{} sin(u)cos(u)du\]

OpenStudy (johnweldon1993):

Now make another substitution.... s = sin(u) ds = cos(u)du \[\large \int^{}_{} s \space ds\]

OpenStudy (johnweldon1993):

Integral of s is \[\large \frac{s^2}{2} + C\]

OpenStudy (johnweldon1993):

Now re-substitute in the original values... s = sin(u) \[\large \frac{sin(u)^2}{2} + C\]

OpenStudy (johnweldon1993):

And u = 2x so \[\large \frac{1}{2}\sin^2(2x) + C\]

OpenStudy (johnweldon1993):

Which then simplifies down to \[\large \frac{1}{4}sin(4x) + C\]

OpenStudy (anonymous):

Wait. where did the 1/2 come from after you plug in u?

OpenStudy (anonymous):

It doesn't have the du for it so... why?

OpenStudy (johnweldon1993):

\(\huge \frac{sin^2(u)}{2} = \frac{1}{2}sin^2(u)\) Just decided to move the 1/2 out in front instead of dividing the whole thing by 2

OpenStudy (anonymous):

Ah... Ok, that makes perfect sense then, sorry. Guess I gotta look up the sin^2(2x) identity later... Would my method of sin(4x) have worked?? Because finding the integral of it would be... u method again? u = 4x uh... would it have?

OpenStudy (johnweldon1993):

Indeed ...actually looking back.you would have received the right answer that way too... \[\large \int^{}_{}sin(4x)dx = -\frac{1}{4}cos(u) + C= -\frac{1}{4}cos(4x) + C\] same difference there

OpenStudy (anonymous):

wouldn't that be wrong though? The answer from sin 4x is cos where as the answer you derived using S is sin... what's going on!?

OpenStudy (anonymous):

still best response, thanks.